40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 14 07:51:20 2011, steve, Update, SUS, PRM damping restored 
    Reply  Thu Sep 15 07:28:08 2011, steve, Update, SUS, PRM damping restored 
Message ID: 5403     Entry time: Wed Sep 14 07:51:20 2011     Reply to this: 5415
Author: steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

The PRM damping was restored at side sensor var 1050

ELOG V3.1.3-